OpenStudy (anonymous):

If \frac{x^2}{25} + \frac {y^2}{64} = 1 and y( 1 ) = 7.83837, find y'( 1 ) by implicit differentiation

OpenStudy (lgbasallote):

\(\Huge \frac{x^2}{25} + \frac {y^2}{64} = 1\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats right....

OpenStudy (anonymous):

any ideas on how to solve??

OpenStudy (lgbasallote):

sadly no :/ i thought i did lol...i just corrected the latex for the better ones to see

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

:(